The path of a ball projected in the air can be represented by the equation h=-16t^2+64t where h represents height, and t represents time. At what times is the ball 48 feet high?
a.)-1 second and -3 seconds
b.)1 second and 3 seconds
c.)1 second and 4 sesconds
d.)1 second and 2 seconds

Found 2 solutions by AAfter Search, juicystarr17:
Answer by AAfter Search(61)

At height 48 feet, -16t^2 + 64t = 48
=> 16t^2 - 64t + 48 = 0
=> t^2 - 4t + 3 = 0
=> t^2 - t - 3t + 3 = 0
=> t(t - 1) - 3(t - 1) = 0
=> (t - 3)(t - 1) = 0
=> t = 3, 1
Hence, the ball is at a height of 48 feet at time 1 sec and 3 sec.
Answer: b

Answer by juicystarr17(1)

At height 48 feet, -16t^2 + 64t = 48
=> 16t^2 - 64t + 48 = 0
=> t^2 - 4t + 3 = 0
=> t^2 - t - 3t + 3 = 0
=> t(t - 1) - 3(t - 1) = 0
=> (t - 3)(t - 1) = 0
=> t = 3, 1
Hence, the ball is at a height of 48 feet at time 1 sec and 3 sec.
Answer: b